OpenStudy (anonymous):

John is choosing a password for his access to the internet. He decided not to use the digit 0 or the letter M. Each letter or number may be used more than once. How many passwords of 2 letters followed by 4 digits are possible?--- is there a formula for this? or a way to plug it into the calculator?

OpenStudy (cwrw238):

two letter part: there are 25 letters used total possibilities for this = no. of permutations of 2 out of 25 = 25*24

OpenStudy (cwrw238):

ooops - the letters can be used more than once so we must add 25 to the above

OpenStudy (anonymous):

it is the counting principle

OpenStudy (cwrw238):

the above is a quick way of doing it the formula is number of permutations r from n = n! / (n-r)!

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

25 letters, 9 digits, \(25\times 25\times 9\times 9\times 9\)

OpenStudy (anonymous):

The choice for the no. of ways in which the first two letters can be chosen from available 25 letters = 25P2 = 600 And, the no. of ways in which the four digits can be chosen from available nine other digits = 9P2=72 Now, The no. of passwords possible are = 600 * 72 = 43200 GOOD LUCK

OpenStudy (anonymous):

not permutations, as you are allowed repetition

OpenStudy (cwrw238):

right - i realized that after my first post

OpenStudy (cwrw238):

@satellite - isn't it 9*9*9*9 ?

OpenStudy (anonymous):

yeah i forgot a 9

OpenStudy (anonymous):

\[25\times 25\times 9\times 9\times 9\times 9\] more succinctly written as \(25^2\times 9^4\)